# Class Distribution Shifts in Zero-Shot Learning: Learning Robust Representations

### v3

@@ Abstract
Zero-shot learning methods typically assume that the new, unseen classes that are encountered at deployment, come from the same distribution as training classes. However, real-world scenarios often involve class distribution shifts (e.g., in age or gender for person identification), posing challenges for zero-shot classifiers that rely on learned representations from training classes. In this work, we propose a model that assumes that the attribute responsible for the shift is unknown in advance, and show that standard training may lead to non-robust representations. To mitigate this, we propose an algorithm for learning robust representations by (a) constructing synthetic data environments via hierarchical sampling and (b) applying environment balancing penalization, inspired by out-of-distribution problems. We show that our approach improves generalization on diverse class distributions in both simulations and real-world datasets.
@@ 1 Introduction
Zero-shot learning [14, 27] tackles classification in scenarios where, in deployment, the classifier encounters new classes that were unseen during training (open-world classification). Zero-shot systems are used in extreme multi-class applications such as face or voice recognition [19] to match multiple observations from the same person, and generally for learning data representations [2].
@@
Class distribution shifts usually refer to changes in class prevalence between training and testing, assuming a fixed set of classes. However, in zero-shot learning, a key question arises: do the new classes follow the same distribution as the training classes? Most zero-shot methods rely on the assumption that the classes are indeed sampled from the same distribution. This assumption is implicitly incorporated into the design of test sets [57, 16], and is explicitly used to assess the ability of zero-shot classifiers to generalize [59, 48].
@@
In practice, training classes are often chosen based on convenience and accessibility during the data collection. Even with careful collection, the distribution of classes may shift over time, leading to a different distribution. For instance, consider a face recognition system deployed in a building located in a neighborhood undergoing demographic changes.
@@
Class distribution shifts pose significant challenges for zero-shot classifiers, since they rely on learning from the training classes to distinguish new, unseen ones. Typically, these classifiers are trained by minimizing the loss over the training set to achieve effective separation among the training classes. However, this approach may result in poor performance when confronted with distributions that significantly differ from the class distribution in the training data. Notably, in person re-identification, this concern gained attention from a fairness perspective with respect to gender bias [15, 23], age bias [5, 33, 50] and racial bias [39, 54]. In all these examples it is known which variable (i.e., gender, age, race) is expected to cause the distribution shift.
@@
However, in real-world scenarios, it is usually unknown during training which attribute will be responsible for a future distribution shift. In such cases, approaches based on collecting balanced datasets or re-weighting training examples [54, 41, 53] are inapplicable. Furthermore, while class distribution shifts have been extensively studied in the standard setting of supervised learning (see Appendix A), this previous work assumes a closed-world setting. It does not account for new classes at test time, and only addresses changes in the prevalence of fixed classes between training and testing. Consequently, class distribution shifts in zero-shot learning remain untreated.
@@
As a first contribution of this work (§3), we study the effects of class distribution shifts on contrastive zero-shot learning, by proposing and analyzing a parametric model. We identify conditions where minimizing loss in this model leads to representations that perform poorly on shifted distributions.
@@
We then use the insights gained from this model to present our second contribution (§4): a first algorithm for learning robust representations against class distribution shifts in zero-shot classifiers. Our proposed approach constructs artificial data environments with diverse attribute distributions using hierarchical sub-sampling, and applies an environment balancing criterion inspired by out-of-distribution (OOD) methods. We assess the effectiveness of our method through both simulations and experiments on real-world datasets, demonstrating its enhanced robustness in §5.
@@ 1.1 Problem Setup
Let $\{z_{i},c_{i}\}_{i=1}^{N_{z}}$ be a labeled set of training data points $z\in\mathcal{Z}$ and classes $c\in\mathcal{C}$, such that $c_{i}$ is the class of $z_{i}$. In this work we focus on verification algorithms that in order to allow for the open-world classification, determine whether a pair of data-points $x_{ij}\coloneqq(z_{i},z_{j})$ belong to the same class. For instance, in person re-identification the task is to identify whether two data points (e.g., face images or voice recordings) belong to the same person. We denote this by $y_{ij}$, where $y_{ij}=1$ if $c_{i}=c_{j}$ and $y_{ij}=0$ otherwise. When the identity of each data point in the pair is not important, we use a single index for simplicity, namely $(x_{k},y_{k})$.
We assume that each class $c$ is characterized by some attribute $A$. We assume that the training classes are sampled from $P_{C}(c)$, the test classes are sampled according to $Q_{C}(c)$, and that the two distributions differ solely due to a shift in the attribute $A$:
@@
We assume that the variable $A$ is unknown, and that both in training and in testing, data points $z\in\mathcal{Z}$ for each class are sampled according to $P_{Z|C}(z|c)$. For instance, revisit the person identification example where each person is a class. If the attribute $A$ is binary (e.g., $a_{1}$ is blond and $a_{2}$ is dark-haired), then $P(C|A=a_{1})$ represents the distribution of blond individuals, and $P(C|A=a_{2})$ over dark haired ones. The training classes might be predominantly sampled from the blond population $P(A=a_{1})=\rho_{\text{tr}}=0.8$, while test classes from $Q(A=a_{1})=\rho_{\text{tr}}=0.1$.
We focus on verification techniques that are based on deep metric learning methods (for reviews see [43, 34]) such as contrastive-learning [17], Siamese neural networks [24], triplet networks [20], and more recent variations [35, 49, 56, 58]. These methods learn a representation function that maps data-points to a representation space $g:\mathcal{Z}\rightarrow\hat{\mathcal{Z}}$, so that examples from the same class are close (in a predefined distance function $d(\cdot,\cdot)$), whereas those from different classes are farther apart.
We assume that $g$ is a neural network that is trained by optimizing a deep-metric-learning loss, such as the contrastive loss [17]:
@@
where $m\geq 0$ is a predefined margin. In our theoretical results, we analyze the no-Hinge contrastive loss (see Appendix B for additional details):
@@
To assess the quality of the representation we treat distances between representations $d_{g}(z_{i},z_{j})\coloneqq d(g(z_{i}),g(z_{j}))$ as classification scores. Following common practice in the field (see, for example, [47, 22]), we use the area under the receiver operating characteristic curve (AUC) to evaluate the representation, allowing for threshold-agnostic assessment:
@@
Our goal is to learn a representation $g$ that is robust to class attribute shifts. That is, such that for an unknown shifted distribution $Q_{A}$, the performance $\mathbb{E}_{Q_{A}}\left[\text{AUC}(g)\right]$ does not significantly deteriorate compared the the performance obtained on the training distribution $P_{A}$.
@@ 2 Background on Environment Balancing Methods in OOD Generalization
The field of OOD generalization gained attention since the work of Peters et al. [36], [37]. It deals with closed-world classification where trainng data is gathered from multiple environments $E_{\text{train}}$. In this setting it is assumed that in each environment $e\in E_{\text{train}}$, examples share the same joint distribution $P^{e}_{C,Z}(c,z)$, but across environments $e$, the joint distribution changes, often due to variations in $P^{e}_{Z|C}(z|c)$. A prominent example [1] involves classifying images of cows and camels and demonstrates how an algorithm relying on background cues during training (e.g., cows in green pastures, camels in deserts) performs poorly on new images of cows in sandy backgrounds.
@@
Most relevant to our work are methods that aim to balance the loss over multiple environments. Consider a representation $g=g_{\theta}$ that is a neural network parameterized by $\theta$, trained to optimize an objective of the form
@@
where $\ell^{e}(g_{\theta})$ is the empirical loss obtained on the environment $e$, $E_{\text{train}}$ is the set of all training environments, $R$ is a regularization term designed to balance performance over multiple environments, and $\lambda$ is a regularization factor weighing between the Empirical Risk Minimization (ERM) term and the balance penalty. We present below three such methods, which we later refer to.
@@
[1] aims to find data representations $g_{\theta}$ such that the optimal classifier $w$ on top of that data representation $w\circ g_{\theta}$, is shared for all environments. The authors therefore propose minimizing the sum of environment losses $\ell^{e}(w\circ g_{\theta})$ over all training environments, such that $w\in\arg\min_{w^{\prime}}\ell^{e}(w^{\prime}\circ g_{\theta})\quad\forall e\in E_{\text{train}}$. However, since it is too hard to optimize, a relaxed objective is proposed instead. It takes the form of Equation 5 with a regularization that measures how close $w$ is to minimizing $\ell^{e}(w\circ g_{\theta})$: $R^{e}_{\text{IRMv1}}(g_{\theta})=\left\|\nabla_{w|w=1}\ell^{e}\left(w\cdot g_{\theta}\right)\right\|^{2}$.
Note that for loss functions, for which optimal classifiers can be written as conditional expectations, the original IRM objective is equivalent to the requirement that for all environments $e,e^{\prime}\in E_{\text{train}}$, $\mathbb{E}_{P_{C,Z}^{e}}\left[c|g(z)=h\right]=\mathbb{E}_{P_{C,Z}^{e^{\prime}}}\left[c|g(z)=h\right],$ where $P_{C,Z}^{e}$ and $P_{C,Z}^{e^{\prime}}$ are the joint data distribution in the respective environments.
@@
[52] leverages the equivalence above to establish a link between multi-environment calibration and invariance for binary predictors ($c\in\{0,1\}$). The proposed regularizer is based on Maximum Mean Calibration Error (MMCE) [26]. Let $s:\hat{\mathcal{Z}}\rightarrow[0,1]$ be a classification score function applied on the representation $s\circ g$, and $s_{i}=\max\{s\circ g(z_{i}),1-s\circ g(z_{i})\}$ be the confidence on the $i$-th data point. Denote by $b_{i}=\mathbbm{1}\{\left|c_{i}-s_{i}\right|<\frac{1}{2}\}$ the correctness, and let $K:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ be a universal kernel. Let $Z^{e}$ denote the training data in the environment $e$. The authors propose using MMCE as a regularization in an objective formed as Equation 5 with $R_{\text{MMCE}}^{e}\left(s,g_{\theta}\right)=\frac{1}{m^{2}}\!\!\sum_{z_{i},z_{j}\in Z^{e}}\!\!\left(b_{i}-s_{i}\right)\left(b_{j}-s_{j}\right)K\left(s_{i},s_{j}\right).$
@@
While simple and intuitive, this approach assumes that losses on different environments accurately reflect the classifier’s performance. However, as discussed in §4, this is often not the case for deep metric learning losses.
@@ 3 Parametric model of class distribution shifts in zero-shot learning
In this section, we introduce a parametric model of class distribution shifts. This model shows that in zero-shot learning, even if the conditional distribution of data given the class $P(z|c)$ remains the same between training and testing, a shift in the class distribution from $P(c)$ to $Q(c)$ can cause poor performance on newly encountered classes, sampled from the shifted distribution $Q(c)$.
Assume that for all classes the data points $z_{i}\in\mathbb{R}^{d}$ are sampled from $z_{i}|c_{i}\sim\mathcal{N}\left(c_{i},\Sigma_{z}\right)$, where $\Sigma_{z}=\nu_{z}\cdot I_{d}$ and $I_{d}$ is the identity matrix. Let the attribute $A$ indicate the type to which a class $c$ belongs, and assume there are two types: $a_{1}$ and $a_{2}$, such that in training $a_{1}$ is the majority type in training with $P(a_{1})=1-\rho_{\text{tr}}\gg 0.5$, while at test time $a_{2}$ is the majority type with $Q(a_{1})=1-\rho_{\text{te}}\ll 0.5$.
Assume that the classes $c_{i}$ are drawn according to $c_{i}\sim\mathcal{N}\left(0,\Sigma_{a}\right)$ for $a\in\{a_{1},a_{2}\}$. We construct the model such that differences between the training class distribution $P(c)$ and the test distribution $Q(c)$ arise solely from a shift in the mixing probabilities of an unknown attribute $A$ (see Equation 1). Therefore, we define $\Sigma_{a}$ as a diagonal matrix with replicates of three distinct values on its diagonal: $\nu_{0},\nu^{+},\nu^{-}$. Let $0<\nu^{-}<\nu_{z}\leq\nu_{0}<\nu^{+}$. Then, in the coordinates corresponding to $\nu_{0}$ and $\nu^{+}$ data points from different classes are well separated, whereas in the coordinates corresponding to $\nu^{-}$ they are not. Assume the coordinates corresponding to $\nu_{0}$ are shared for both types, but $\nu^{+}$ and $\nu^{-}$ are swapped:
@@
Figure 1: Illustration of the parametric model. Data points $z$ are shown in three dimensions, with type $a_{1}$ represented by red triangles and type $a_{2}$ by green circles. Each class is best separated along specific axes: $a_{1}$ along the red ($z^{(1)}$) and $a_{2}$ along the green ($z^{(2)}$). On the axis $z^{(0)}$ both types can be separated, but not as effectively as on their respective optimal axes.
Figure 2: Optimal weight values for varying variance and dimension ratios, as a function of $\rho$.
The following proposition shows that if the number of components $d_{2}$ corresponding to $\nu^{-}$ for the training classes is not too large, that is $d_{2}<h\left(\rho,\nu_{z},\nu_{0},\nu_{1},\nu_{2}d_{1}\right)$, the optimal solution for the training distribution prioritizes the components (features) corresponding to $\nu^{+}$ for classes of type $a_{1}$. Note that if $d_{2}$ is large, when combined, the corresponding components may still provide reasonable separation. We define $h$ in equation 35, and provide the proof of Proposition 1 in Appendix B.2.
@@
Consider a weight representation $g(z)=Wz$, where $W\in\mathbb{R}^{d\times d}$is a diagonal matrix $W=\text{diag}(w)$, and the squared Euclidean distance $d_{g}\left(z_{i},z_{j}\right)=\left\|W\left(z_{i}-z_{j}\right)\right\|^{2}$. Let $W^{*}=\text{diag}(w^{*})$ be such that $W^{*}\in\arg\min_{W}\mathbb{E}\left[\widetilde{\ell}\left(\cdot,\cdot,\cdot;d_{g}\right)\right]$. Denote $\overline{w}^{*2}_{1}=\frac{1}{d_{1}}\sum_{k=d_{0}+1}^{d_{1}}w^{*}_{k}$ and $\overline{w}^{*2}_{2}=\frac{1}{d_{2}}\sum_{k=d_{1}+1}^{d}w^{*2}_{k}$. Then for all $\rho>\frac{1}{2}$, and $\nu_{z},\nu_{0},\nu_{1},\nu_{2},d_{1},d_{2}$ satisfying $d_{2}<h\left(\rho,\nu_{z},\nu_{0},\nu_{1},\nu_{2},d_{1}\right)$ we have $d_{2}\overline{w}^{*2}_{2}\leq d_{1}\overline{w}^{*2}_{1}$.
Since components corresponding to $\nu^{+}$ for classes of type $a_{1}$ align with $\nu^{-}$ for classes of type $a_{2}$, the optimal representation on the training distribution yields poor separation on the shifted test distribution. Therefore, a robust representation should prioritize dimensions that provide effective separation for both class types, corresponding to $\nu_{0}$.
In Appendix B.3 we provide the full analytical solution $w^{*}$ that minimizes the expected loss $\mathbb{E}\left[\widetilde{\ell}\left(\cdot,\cdot,\cdot;d_{g}\right)\right]$ or the weight representation $g(z)=Wz$, using the squared Euclidean distance. According to Proposition 1, larger $d_{2}$ values favor $\nu^{-}$ for better aggregated separation. Increasing $\nu_{0}/\nu^{+}$ leads to growing differences between ${w}^{*2}_{1}$ and ${w}^{*2}_{2}$, and vice versa for $\nu_{0}/\nu^{-}$.
This aligns with a common consideration in the Out-of-Distribution (OOD) generalization field (see §2), often refereed to as invariance.

### v4

@@ Abstract
Zero-shot learning methods typically assume that the new, unseen classes encountered during deployment come from the same distribution as the the classes in the training set. However, real-world scenarios often involve class distribution shifts (e.g., in age or gender for person identification), posing challenges for zero-shot classifiers that rely on learned representations from training classes. In this work, we propose and analyze a model that assumes that the attribute responsible for the shift is unknown in advance. We show that in this setting, standard training may lead to non-robust representations. To mitigate this, we develop an algorithm for learning robust representations in which (a) synthetic data environments are constructed via hierarchical sampling, and (b) environment balancing penalization, inspired by out-of-distribution problems, is applied. We show that our algorithm improves generalization to diverse class distributions in both simulations and experiments on real-world datasets.
@@ 1 Introduction
Zero-shot learning systems [14, 27] are designed to classify instances of new, previously unseen classes at deployment, a scenario known as open-world classification. These systems are widely applied in extreme multi-class applications, such as face or voice recognition [19] for matching observations of the same individual, and more generally, for learning data representations [2].
@@
Class distribution shifts typically refer to changes in the prevalence of a fixed set of classes between training and testing. In zero-shot learning, however, a different challenge arises: the appearance of entirely new classes at test time. This raises a critical question – are these new classes drawn from the same distribution as the training classes? Most zero-shot methods assume that they are, an assumption that not only shapes the design of test sets [57, 16] but also plays an explicit role in assessing the generalization capabilities of zero-shot classifiers [59, 48].
@@
In practice, training classes are often chosen based on convenience and accessibility during data collection. Even when data is carefully collected, the distribution of classes may shift over time, leading to a different distribution. For instance, this could occur when a face recognition system is deployed in a building located in a neighborhood undergoing demographic changes.
@@
Class distribution shifts pose significant challenges to zero-shot classifiers, since they rely on learning data representations from the training classes to distinguish new, unseen ones. Typically, these classifiers are trained by minimizing the loss on the training set to effectively separate the training classes. However, this approach may result in poor performance when confronted with data from distributions that differ significantly from the class distribution in the training data. Notably, in person re-identification, this concern gained attention from a fairness perspective with respect to gender [15, 23], age [5, 33, 50], and racial [39, 54] bias. In all these studies the variable (i.e., gender, age, race) expected to cause the distribution shift was known in advance.
@@
In contrast, in real-world scenarios, the attribute responsible for a future distribution shift is usually unknown during training. In such cases, existing approaches based on collecting balanced datasets or re-weighting training examples [54, 41, 53] are inapplicable. Furthermore, while class distribution shifts have been extensively studied in the standard setting of supervised learning (see Appendix A), previous research assumed a closed-world setting that does not account for new classes at test time. Instead, it only addressed changes in the prevalence of fixed classes between training and testing. Consequently, class distribution shifts in zero-shot learning remain largely unaddressed.
@@
In this paper we first address these limitations by examining the effects of class distribution shifts on contrastive zero-shot learning, by proposing and analyzing a parametric model (§3). We identify conditions where minimizing loss in this model leads to representations that perform poorly when a distribution shift has occurred.
@@
We then use the insights gained from this model to present our second contribution (§4): an algorithm for learning representations that are robust against class distribution shifts in zero-shot classification. In our proposed approach, artificial data environments with diverse attribute distributions are constructed using hierarchical subsampling, and an environment balancing criterion inspired by out-of-distribution (OOD) methods is applied. We assess our method’s effectiveness in both simulations and experiments on real-world datasets, demonstrating its enhanced robustness in §5.
@@ 1.1 Problem Setup
Let $\{z_{i},c_{i}\}_{i=1}^{N_{z}}$ be a labeled set of training data points $z\in\mathcal{Z}$ and classes $c\in\mathcal{C}$, such that $c_{i}$ is the class of $z_{i}$.
In this work, we focus on verification algorithms that enable open-world classification by determining whether two data points $x_{ij}\coloneqq(z_{i},z_{j})$ belong to the same class. For instance, in person re-identification the task is to identify whether two data points (e.g., face images or voice recordings) belong to the same person. We denote this by $y_{ij}$, where $y_{ij}=1$ if $c_{i}=c_{j}$ and $y_{ij}=0$ otherwise. When the identity of each data point in the pair is not important, a single index is used for simplicity, namely $(x_{k},y_{k})$.
We assume that each class $c$ is characterized by some attribute $A$. We further assume that the training classes are sampled from $P_{C}(c)$, the test classes are sampled according to $Q_{C}(c)$, and the two distributions differ solely due to a shift in the distribution of an attribute $A$:
@@
Importantly, we assume that the attribute $A$ is unknown, and that both during training and testing, data points $z\in\mathcal{Z}$ for each class are sampled according to $P_{Z|C}(z|c)$. For instance, revisit the person identification example where each person is a class. If the attribute $A$ is binary (e.g., $a_{1}$ is blond and $a_{2}$ is dark-haired), then $P(C|A=a_{1})$ represents the distribution of people with blond hair, and $P(C|A=a_{2})$ of individuals with other hair colors. The training classes might be predominantly sampled from the blond population $P(A=a_{1})=\rho_{\text{tr}}=0.8$, while test classes are predominantly sampled from $Q(A=a_{1})=\rho_{\text{te}}=0.1$.
We focus on verification techniques based on deep metric learning methods (for surveys see [43, 34]) such as contrastive-learning [17], Siamese neural networks [24], triplet networks [20], and other more recent variations [35, 49, 56, 58]. These methods learn a representation function that maps data points to a representation space $g:\mathcal{Z}\rightarrow\hat{\mathcal{Z}}$, so that examples from the same class are close (in a predefined distance function $d(\cdot,\cdot)$), while those from different classes are farther apart.
We assume that $g$ is a neural network trained by optimizing a deep-metric-learning loss, such as the contrastive loss [17]:
@@
where $m\geq 0$ is a predefined margin, and $d_{g}(z_{i},z_{j})\coloneqq d(g(z_{i}),g(z_{j}))$ is the distance between the representations of the datapoints $z_{i},z_{j}$. In our theoretical analysis, we examine the no-hinge contrastive loss (see Appendix B for additional details):
@@
To evaluate the class separation capability of a representation, we treat the distances between representations, $d_{g}(z_{i},z_{j})$, as classification scores. Following common practice in the field (e.g., [47, 22]), we use the area under the receiver operating characteristic curve (AUC) to evaluate the representation, enabling threshold-agnostic assessment:
@@
Our goal is to learn a representation $g$ that is robust to class attribute shifts. That is, such that for an unknown shifted distribution $Q_{A}$, the performance $\mathbb{E}_{Q_{A}}\left[\text{AUC}(g)\right]$ does not significantly deteriorate compared to the performance obtained on the training distribution $P_{A}$.
@@ 2 Background on Environment Balancing Methods in OOD Generalization
The field of OOD generalization gained attention since the work of Peters et al. [36], [37], which deals with closed-world classification where training data is gathered from multiple environments $E_{\text{train}}$. In this setting it is assumed that in each environment $e\in E_{\text{train}}$ examples share the same joint distribution $P^{e}_{C,Z}(c,z)$, but across environments the joint distribution changes, often due to variations in $P^{e}_{Z|C}(z|c)$. A well-known example [1] involving the classification of images of cows and camels demonstrates how an algorithm relying on background cues during training (e.g., cows in green pastures, camels in deserts) performs poorly on new images of cows with sandy backgrounds.
@@
Most relevant to our work are methods that aim to balance the loss over multiple environments. These methods consider a representation $g=g_{\theta}$ that is a neural network parameterized by $\theta$ trained to optimize an objective of the form
@@
where $\ell^{e}(g_{\theta})$ is the empirical loss obtained on the environment $e$, $E_{\text{train}}$ is the set of all training environments, $R$ is a regularization term designed to balance performance over multiple environments, and $\lambda$ is a regularization factor balancing the tradeoff between the empirical risk minimization (ERM) term and the balance penalty. Below, we describe three such methods, which we will refer to later in the paper.
@@
presented in [1], aims to find data representations $g_{\theta}$ such that the optimal classifier $w$ on top of the data representation $w\circ g_{\theta}$ is shared across all environments. Therefore, the authors proposed minimizing the sum of environment losses $\ell^{e}(w\circ g_{\theta})$ over all training environments such that $w\in\arg\min_{w^{\prime}}\ell^{e}(w^{\prime}\circ g_{\theta})$ for all $e\in E_{\text{train}}$. However, since this objective is too difficult to optimize, a relaxed version was also proposed, taking the form of Equation 5 with a penalty that measures how close $w$ is to minimizing $\ell^{e}(w\circ g_{\theta})$: $R^{e}_{\text{IRMv1}}(g_{\theta})=\left\|\nabla_{w|w=1}\ell^{e}\left(w\cdot g_{\theta}\right)\right\|^{2}$.
Note that for loss functions for which optimal classifiers can be expressed as conditional expectations, the original IRM objective is equivalent to the requirement that for all environments $e,e^{\prime}\in E_{\text{train}}$, $\mathbb{E}_{P_{C,Z}^{e}}\left[c|g(z)=h\right]=\mathbb{E}_{P_{C,Z}^{e^{\prime}}}\left[c|g(z)=h\right],$ where $P_{C,Z}^{e}$ and $P_{C,Z}^{e^{\prime}}$ are the joint data distributions in the respective environments.
@@
presented in [52], leverages the equivalence above to establish a link between multi-environment calibration and invariance for binary predictors ($c\in\{0,1\}$). The proposed regularizer is based on the maximum mean calibration error (MMCE) [26]. Let $s:\hat{\mathcal{Z}}\rightarrow[0,1]$ be a classification score function applied on the representation $s\circ g$, and $s_{i}=\max\{s\circ g(z_{i}),1-s\circ g(z_{i})\}$ be the confidence on the $i$-th data point. Denote the correctness as $b_{i}=\mathbbm{1}\{\left|c_{i}-s_{i}\right|<\frac{1}{2}\}$, and let $K:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ be a universal kernel. Let $Z^{e}$ denote the training data in the environment $e$. The authors proposed using the MMCE as a penalty in an objective that takes the form of Equation 5 with $R_{\text{MMCE}}^{e}\left(s,g_{\theta}\right)=\frac{1}{m^{2}}\!\!\sum_{z_{i},z_{j}\in Z^{e}}\!\!\left(b_{i}-s_{i}\right)\left(b_{j}-s_{j}\right)K\left(s_{i},s_{j}\right).$
@@
While simple and intuitive, this approach assumes that losses across different environments accurately reflect the classifier’s performance. However, as discussed in §4, this is often not true for deep metric learning losses, where significant changes in loss may correspond to only minor variations in performance.
@@ 3 Parametric Model of Class Distribution Shifts in Zero-Shot Learning
In this section, we introduce a parametric model of class distribution shifts. Our model shows that in zero-shot learning, even if the conditional distribution of data given the class $P(z|c)$ remains the same between training and testing, a shift in the class distribution from $P(c)$ to $Q(c)$ can cause poor performance on newly encountered classes sampled from the shifted distribution $Q(c)$.
Assume that for all classes, the data points $z_{i}\in\mathbb{R}^{d}$ are sampled from $z_{i}|c_{i}\sim\mathcal{N}\left(c_{i},\Sigma_{z}\right)$, where $\Sigma_{z}=\nu_{z}\cdot I_{d}$, and $I_{d}$ is the identity matrix. Let the attribute $A$ indicate the type of a class $c$, with two possible types: $a_{1}$ and $a_{2}$. Assume that the classes $c_{i}$ are drawn according to $c_{i}\sim\mathcal{N}\left(0,\Sigma_{a}\right)$ for $a\in\{a_{1},a_{2}\}$. Finally, assume that in training, $a_{1}$ is the majority type with $P(a_{1})=\rho_{\text{tr}}\gg 0.5$, whereas at test time, $a_{2}$ is the majority type with $Q(a_{1})=\rho_{\text{te}}\ll 0.5$.
We construct the model such that differences between the training class distribution $P(c)$ and the test distribution $Q(c)$ stem solely from a shift in the mixing probabilities of an unknown attribute $A$ (see Equation 1). Therefore, we define $\Sigma_{a}$ as a diagonal matrix with replicates of three distinct values on its diagonal: $\nu_{0},\nu^{+},\nu^{-}$. Let $0<\nu^{-}<\nu_{z}\leq\nu_{0}<\nu^{+}$. Then, in the coordinates corresponding to $\nu_{0}$ and $\nu^{+}$ data points from different classes are well separated, whereas in the coordinates corresponding to $\nu^{-}$ they are not. Assume the coordinates corresponding to $\nu_{0}$ are shared by both types, but $\nu^{+}$ and $\nu^{-}$ are swapped:
@@
Figure 1: Illustration of the parametric model. Classes of each type are best separated along specific axes: classes of type $a_{1}$ along the red axis ($z^{(1)}$) and classes of type $a_{2}$ along the green axis ($z^{(2)}$). On axis $z^{(0)}$ both types can be separated but not as effectively as on their respective optimal axes.
Figure 2: Optimal weights. Top row: $d_{0}$ is fixed, $d_{1}$ and $d_{2}$ vary. Middle and bottom rows: $d_{0},d_{1},d_{2}$ are fixed. Middle: $\nicefrac{{\nu_{0}}}{{\nu^{-}}}$ varies. Bottom: $\nicefrac{{\nu_{0}}}{{\nu^{+}}}$ varies.
The following proposition shows that if the number of dimensions $d_{1}$ that allow good separation for classes of type $a_{1}$ is relatively similar to the number of dimensions $d_{2}$ that enable good separation for classes of type $a_{2}$, specifically if $h_{l}\left(\rho,\nu_{z},\nu_{0},\nu_{1},\nu_{2}\right)<\frac{d_{2}+2}{d_{1}+2}<h_{u}\left(\rho,\nu_{z},\nu_{0},\nu_{1},\nu_{2}\right)$, then the optimal solution for the training distribution prioritizes the components (features) corresponding to $\nu^{+}$ for classes of type $a_{1}$. Thus, the prioritized features allow good separation for classes from the majority type in training, but offer poor separation for the shifted test distribution, where most classes are of type $a_{2}$. Note that if $d_{2}$ is large, when combined, the corresponding components may still provide reasonable separation. We define $h_{l}$ and $h_{u}$ in Equation 35 and provide the proof of Proposition 1 in Appendix B.2.
@@
Consider a weight representation $g(z)=Wz$, where $W\in\mathbb{R}^{d\times d}$is a diagonal matrix, and the squared Euclidean distance $d_{g}\left(z_{i},z_{j}\right)=\left\|W\left(z_{i}-z_{j}\right)\right\|^{2}$. Let $W^{*}=\text{diag}(w^{*})\in\arg\min_{W}\mathbb{E}\left[\widetilde{\ell}\left(\cdot,\cdot,\cdot;d_{g}\right)\right]$. Denote $\overline{w}^{*2}_{1}=\frac{1}{d_{1}}\sum_{k=d_{0}+1}^{d_{1}}w^{*}_{k}$ and $\overline{w}^{*2}_{2}=\frac{1}{d_{2}}\sum_{k=d_{1}+1}^{d}w^{*2}_{k}$. Then, for all $\rho>\frac{1}{2}$ and $\nu_{z},\nu_{0},\nu_{1},\nu_{2},d_{1},d_{2}$ satisfying $h_{l}\left(\rho,\nu_{z},\nu_{0},\nu_{1},\nu_{2}\right)<\frac{d_{2}+2}{d_{1}+2}<h_{u}\left(\rho,\nu_{z},\nu_{0},\nu_{1},\nu_{2}\right)$ it holds that $d_{2}\overline{w}^{*2}_{2}\leq d_{1}\overline{w}^{*2}_{1}$.
Note that the conditions outlined in Proposition 1 are sufficient but not necessary. Accordingly, in Appendix B.3, we provide the complete analytical solution for $w^{*}$ that minimizes the expected loss $\mathbb{E}\left[\widetilde{\ell}\left(\cdot,\cdot,\cdot;d_{g}\right)\right]$ for the weight representation $g(z)=Wz$, using the squared Euclidean distance. According to Proposition 1, larger $d_{2}$ values favor $\nu^{-}$ for better aggregated separation. Increasing $\nu_{0}/\nu^{+}$ leads to increased differences between ${w}^{*2}_{1}$ and ${w}^{*2}_{2}$, and vice versa for $\nu_{0}/\nu^{-}$.
These relationships in the optimal solution are illustrated in Figure 2, showcasing different scenarios. The top row shows that when $d_{1}=d_{2}=10$ dimensions favoring classes of type $a_{1}$ are prioritized for $\rho>0.5$, while those favoring type $a_{2}$ are prioritized for $\rho<0.5$. When $d_{1}=10$ while $d_{2}=5$, dimensions favoring type $a_{1}$ are prioritized for all values of $\rho$, and vice versa when $d_{2}$ is significantly larger than $d_{1}$. The middle and the bottom row further explore the $d_{1}=d_{2}$ case, showing how differences in separability between shared dimensions ($\nu_{0}$) and type-favoring dimensions impact weight allocation.
Since components corresponding to $\nu^{+}$ for classes of type $a_{1}$ align with $\nu^{-}$ for classes of type $a_{2}$, the optimal representation for the training distribution results in poor separation for the shifted test distribution. Therefore, a robust representation should prioritize dimensions that provide effective separation for both class types, corresponding to $\nu_{0}$.
This aligns with a common principle in the OOD generalization field, where robust representations are those that rely on features shared across environments (see §2). This principle is often referred to as invariance.